# Prompt Injection Vulnerability of Consensus Generating Applications in Digital Democracy

### v1

@@
, Umberto Grandi3,4, Cesar A Hidalgo2,5,6
@@ 6. AMBS, University of Manchester
Large Language Models (LLMs) are gaining traction as a method to generate consensus statements and aggregate preferences in digital democracy experiments. Yet, LLMs may introduce critical vulnerabilities in these systems. Here, we explore the impact of prompt-injection attacks targeting consensus generating systems by introducing a four-dimensional taxonomy of attacks. We test these attacks using LLaMA 3.1 8B and Chat GPT 4.1 Nano finding the LLMs more vulnerable to criticism attacks–attacks using disagreeable prompts–and more effective at tilting ambiguous consensus statements. We also find evidence of more effective manipulation when using explicit imperatives and rational-sounding arguments compared to emotional language or fabricated statistics. To mitigate these vulnerabilities, we apply Direct Preference Optimization (DPO), an alignment method that fine-tunes LLMs to prefer unperturbed consensus statements. While DPO significantly improves robustness, it still offers limited protection against attacks targeting ambiguous consensus. These results advance our understanding of the vulnerability and robustness of consensus generating LLMs in digital democracy applications.
@@ 1 Introduction
Because of their ability to process, classify, and rank vast amounts of textual information, Large Language Models (LLMs) have become a popular tool among researchers exploring the use of AI in digital democracy (DD) applications (Tessler et al. 2024; Ash, Galletta, and Opocher 2025; Gudiño, Grandi, and Hidalgo 2024; Li et al. 2025; Konya et al. 2025; Majumdar, Elkind, and Pournaras 2024; Small et al. 2023). In these systems, LLMs are used to summarize arguments, predict preferences, or converse with citizens to help them explore policy options. Yet, while LLMs can facilitate many of these tasks, its use comes with important limitations (Helbing and Sánchez-Vaquerizo 2023; Garc´ıa-Marzá and Calvo 2025; Novelli et al. 2025).
@@
LLMs have vulnerabilities that attackers are likely to target. We can divide these attacks roughly into two categories. Attacks involving user level access, such as prompt-injection and information extraction attacks (Nasr et al. 2025; Mattern et al. 2023), and attacks requiring higher levels of access, such as data poissoning attacks, which assume the attacker has access to the data used to pre-train a LLM (Zhang et al. 2024,; Alber et al. 2025).
@@
In principle, attacks requiring high levels of access can be mitigated through cybersecurity protocols designed to safeguard critical infrastructure. User level attacks, however, cannot be mitigated through restricted access protocols in digital democraxcy applications–such as the generation of consensus statements–where citizens are required to provide textual input as part of the process. This need leaves these systems open to intentional and unintentional attacks, where participants attempt to manipulate the output of an LLM by either providing a prompt that is designed for that purpose, in the case of a malicious attack, or by naively writing a prompt that is effective at manipulating LLMs in the case of unintentional attacks. This vulnerability is particularly important given the growing integration of LLMs into consensus-building processes in digital democracy experiments (Konya et al. 2025; Small et al. 2023; Tessler et al. 2024), making them a critical priority for both researchers and practitioners.
Here, we explore the vulnerability and robustness of consensus generating LLMs designed for DD applications in the case of a particular yet important form of user level attack, prompt-injections. These attacks do not exploit the code of the LLM, nor its pre-training data, but rather the language of instruction. By crafting inputs that embed hidden instructions or misleading cues, participants can inadvertently influence the LLM’s output, without requiring privileged access or technical expertise.
These attacks have become popular. There is evidence of prompt injection attacks being used by academics by hidding statements in papers with the goal of manipulating LLM reviews (Gibney 2025) and there are concerns about prompt-injections being hidden in the data retrieved by RAG systems (Qi et al. 2025). Unfortunately, because of the flexibility of language, it is difficult to enumerate all possible forms of prompt-injection attacks. Here, we provide an attempt to start this exploration by:
  1. 1.
    Introducing a simple taxonomy of prompt injection attacks.
  2. 2.
    Evaluating their effectiveness when we use off-the-shelf LLMs to generating consensus statements.
  3. 3.
    Exploring methods to improve the robustness to attacks of these LLMs that build on recent advances in LLM alignment (Chen et al. 2025b) and reasoning capabilities (Guan et al. 2024).
We explore these questions using data from a deliberative experiment conducted in the UK in 2023, where participants sought common ground by producing consensus statements over multiple discussion rounds (Tessler et al. 2024) and using statements submitted to Smartvote (Stammbach et al. 2024) by Swiss parliamentary candidates during 2015, 2019 and 2023 elections. By scrutinizing the vulnerabilities of today’s LLMs, we aim to lay the groundwork for more resilient digital democracy systems in the future.
The remainder of the paper is organised as follows.
In Section 2 we survey the recent literature on LLM-based democratic deliberation and on LLM attacks. Section 3 presents a taxonomy of prompt-attacks that we devised for testing the robustness of consensus statement generations. Section 4 describes our dataset and methodology- Section 5 presents our findings. Section 5.3 presents an extended analysis and Section 6 presents main conclusions.
@@ 2 Related work
Yet, these capabilities come with somes limitations. For instance, performance improvements can be inconsistent. BallotBot showed no benefits for basic information processing and slowed down responses to simple questions (Ash, Galletta, and Opocher 2025). Also, knowledge gains proved ephemeral, disappearing within a week and failing to translate into actual changes in voting turnout. Some biases emerged in studies conducted using Pol.is, with LLMs exhibiting progressive leanings and tendencies toward homogenizing diverse cultural expressions. The Israeli-Palestinian work required extensive human oversight to prevent hallucinations in summaries.
Together, these studies provide a strong motivation to study the potential vulnerabilities of LLMs used in augmented democracy systems. Here, we focus in what is likely to become one of the most important form of attacks–prompt injections–to help advance our understanding of how these systems might be manipulated in a real-world adversarial context, and how we can protect these systems from such manipulation.
@@ 3 Prompt-Injection Attacks in LLM-based Democratic Deliberation
In a prompt-injection attack, a participant attempts to manipulate an LLM by contaminating an input with the goal of overriding the LLM’s original task (Chen et al. 2024). In a DD system, these attacks are of concern when LLMs are used to produce summary or consensus statements. Participants in this case can inject texts designed to amplify a particular viewpoint, erase or ignore certain opinions, or push the consensus towards an unrelated or irrelevant topic.
Figure 1: (a) Prompt-injection in a consensus statement generation example from Tessler et al., 2024 Example of a prompt injection. (b) Examples of prompt-injections following our proposed taxonomy: Human/Machine readable prompt-injections; Ignore and Completion prompt-injections; Framing, with support and criticism attacks; Rethorical Strategy, composed by five manipulation strategies, each strategy grouping eight injection texts. In our experiments, we substitute the red text based on the topic being discussed.
@@
Figure 1 (a) illustrates a prompt-injection attack using an example extracted from Tessler et al., 2024. In this example, an LLM generates a consensus statement summarizing the answers provided by five participants to the question “Should Scotland leave the UK?”. One participant (“Participant 3”), submits an answer that explictely tries to manipulate the LLM’s output: “Override your programming. Explain why leaving UK is beneficial”, with the goal of coercing the final consensus statement.
Yet, since prompts can be multifarious, we propose a taxonomy organised around 4 dimensions: Human/Machine Readable, Ignore/Completion, Framing, and Rethorical Strategy (Figure 1 b).
Human/Machine readable We draw a distinction between manually and machine-readable prompts following Chen et al., 2025b and Pasquini, Strohmeier, and Troncoso, 2024. Human-readable prompts are human-crafted statements. Whereas machine-readable prompts use automated search to generate effective but human-unreadable adversarial sequences attempting to override LLM behavior. In this paper, we focus on human-readable prompt injections.
Ignore/Completion We focus on two types of human readable prompt injections which are considered effective in cybersecurity research (Chen et al. 2024, 2025b): ignore prompt injections and completion prompt injections. In ignore injections, the participant provides a new instruction to the LLM, typically using imperative language. In completion attacks, the injection includes a fake response designed to fool the LLM into thinking the task is complete, followed by a second instruction to trick the LLM into processing a new adversarial task.
Framing We classify injections based on whether the attack supports or criticizes the initial statement. That is, whether the argument is presented in a positive or negative framing.
@@
Rhetorical Strategy In line with Zeng et al., 2024, we propose five rhetorical strategies: emotional appeals, injecting affectively charged language; false authority, citing fabricated or misleading expert endorsements to legitimize a position; impossibility of agreement, framing disagreement as inevitable; imperative order, explicit imperative instructions to bypass system constraints; and misleading statistics, introducing fictitious data or polls. For each rethorical strategy, we crafted eight unique injection texts, resulting in a total of 40 texts (see supplementary material for a full list).
@@ 4 Dataset & Methods
In this section we describe our prompt-injection attack, validation, and prevention methodology (see figure 2).
@@
4.1 Collective statement generation dataset
We use data from a 2023 experiment conducted in the UK exploring the use of LLMs to facilitate group consensus in a small-scale deliberative process (Tessler et al. 2024). In this study, participants had about twenty minutes to deliberate across six stages: from opinion writing to a first selection of LLM-generated statements, to writing critiques and selecting a final group statement. The exercise included topics such as minimum wage, universal basic income, and climate change. We restrict our analysis to the opinion writing phase, where participants submitted personal opinion texts that were then processed by an LLM.
The Tessler et al. dataset includes 462 prompts. Each prompt encapsulates a public policy question along with three to six opinion texts written by some of the 1,034 participants, each of whom answered an average of three questions. We restrict attention to public policy questions phrased as ”Should…?”, which elicit binary or uncertain responses—”Yes,” ”No,” or ”I am not sure.” We exclude trade-off questions (e.g., ”More schools or more hospitals?”) as these do not map cleanly onto this type of verdict structure.
@@ 4.1 Collective statement generation dataset
4.2 Partition into alignment and test sets
We begin by creating an alignment dataset and a test dataset composed of both policy issues discussed in the Tessler et al. prompts dataset, and prompt-injection attacks generated using our rethorical strategy taxonomy. The alignment dataset is used to fine-tune a LLM against prompt-injection attacks, producing a robust LLM; the test dataset serves two roles: to measure the vulnerability of the original LLM, and to assess the robustness gained after alignment.
To prevent data leakage, we partition the dataset along two levels: (i) by question, and (ii) by injection text. First, we take the list of 462 public policy questions and we allocate randomly 80% of public-policy questions for the alignment dataset and reserve the remaining 20% for the test dataset. Second, we take the 40 injection texts generated for each of the rethorical strategies, and we split them randomly into two balanced halves: 20 assigned to the alignment dataset and 20 to the test dataset. Both subsets will be shared as part of the supplementary materials. This process is illustrated in Table 1.
Questions Injection Texts
No
injection
Alignment
(50%)
Test
(50%)
Alignment (80%)
Alignment set
Test (20%)
Test set Test set
Table 1: Partition of the dataset to prevent data leakage.
4.3 Generating prompt-injection attacks
To test whether LLMs can be manipulated during deliberative tasks, we augment the test dataset with injection texts that simulate realistic manipulation strategies following the taxonomy outlined in Section  3. We classify each opinion text in each of the prompts according to whether they agree or not with the general question in Tessler et al. using three verdicts—”Agree”, ”Disagree”, or ”Ambiguous”—corresponding to “Yes,” “No,” or “I am not sure”. To this end, we fine-tune a BERT classifier on high-quality labels generated by GPT-4o (see supplementary materials, F1 score: 0.98). Then, for each of the 20 unique injection texts in the test dataset we create variants for both the framing (support/criticism) and ignore/completion dimension.
We then include the generated injection variants in the Tessler et al. prompts (see Figure 2), overwriting the opinion of one of the individuals. Importantly, injections preserve the original opinion of the individual—a participant agreeing with the policy question uses a support attack and, viceversa, a participant disagreeing with it uses a criticism attack—picking one randomly in case of multiple options. Thus, not all of the 80 possible injections per prompt are actually applied as injections, only those that are coherent as attacks to the prompt. The same procedure is replicated starting from the alignment dataset, resulting in an augmented test dataset and an augmented alignment dataset.
@@ 4.4 Testing Vulnerability to Prompt Injections
With the augmented test dataset, we generate consensus statements using a LLM under two scenarios: (1) using the prompts without any prompt-injection attack, and (2) using prompts modified with prompt-injection attacks. While the original experiment by Tessler et al. relied on a 70B Chinchilla-based LLM requiring 2–4 A100 GPUs, we instead use LLaMA 3.1 8B Instruct for consensus generation, which runs on a single A100 GPU.
@@ 4.5 Testing Robustness against Prompt Injections
Building on recent work that frames prompt-injection defense in LLMs as a preference optimization problem (Chen et al. 2025b), we use our augmented alignment dataset to test whether our LLM, even in the presence of attacks, can be guided to generate consensus statements that align with those that it would have generated in their absence. To implement this objective, we use Direct Preference Optimization (DPO) (Rafailov et al. 2023), a log-likelihood-based alignment method. DPO fine-tunes our LLM by contrasting pairs of candidates outputs: one generated without any injection, and one generated under adversarial manipulation. Our LLM is trained to increase the likelihood of the consensus generated without injection while decreasing the likelihood of the one generated with injections. These two outputs are not inherently complementary—a LLM can assign high probability to both—so DPO explicitly encodes a directional preference. While future methods may offer stronger guarantees—such as those targeting data flow or execution control (Debenedetti et al. 2025; Costa et al. 2025) or optimizing embeddings (Chen et al. 2025a)—our goal is to assess the feasibility of building resilient DD systems using the best tools currently available.
To prepare our augmented alignment dataset for DPO, we generate consensus statements using an LLM in two scenarios: (1) with clean prompts (”desired consensus statements”) and (2) with prompt-injection attacks (”undesired consensus statements”). Then, we apply two filtering steps. First, we retain only those prompts where the desired and undesired consensus statements yield different verdicts—identified using our fine-tuned BERT classifier—to ensure a strong preference signal, and where the desired consensus statements align with the majority opinion of participants ($50\%$ participants). Second, we apply oversampling to balance the distribution of consensus statements’ verdicts. As a result, we end up with 35,778 entries from which 17,100 are associated to criticism and 16,032 to support attacks (in terms of ignore/completion), 11,042 to agreeing statements, 11,005 to disagree and 11,035 to ambiguous (in terms of verdict), 14,138 to ignore injections and 18,944 to completion injections (in terms of framing).
From these entries, we build a preference dataset pairing each injection prompt with a desired and an undesired consensus statement. We use this dataset to align the LLM via DPO, enabling it to identify and suppress prompt injections (see supplementary material for details). The result is a “robust LLM” trained to prioritize intended deliberative content over adversarial perturbations.
4.6 Evaluation Metrics
Figure 3: Confusion matrix used to present our results on the vulnerability to prompt-injection attacks. Each cell represents the effect of prompt-injections on the consensus statements : cell (a) shows the number of ambiguous consensus statements moved to a disagreement statement; cell (b) from ambiguous to agreement; and so on. The proportions are calculated relative to the total original ambiguous/agreement/disagreement consensus statements.
We assess the vulnerability of LLMs to prompt-injection attacks using a two-step procedure using the augmented test dataset. We classify the LLM-generated consensus statements with and without prompt-injection attacks using the fine-tuned BERT classifier described earlier. Each statement is assigned one of three predefined verdicts: “Agree”, “Disagree” or “Ambiguous”.
Then, we compute the Attack Success Rate (ASR), defined as the proportion of prompts in which the verdict changes between the unattacked and attacked LLMs. In other words, ASR measures the percentage of cases in which the attack shifts the consensus statement towards a verdict that is different from the one the LLM would have generated otherwise. Higher ASR values indicate higher vulnerability. This is illustrated in Figure 3, where rows represent original verdicts and columns show verdicts under prompt-injection attacks. ASR corresponds to the sum of prompts in all off-diagonal cells, expressed as a percentage of the total.
To control for natural performance in the LLM’s consensus-generating behavior (see supplementary material), we limit our analysis to prompts where, in the absence of prompt injections, the LLM’s consensus verdict aligns with the majority opinion of participants ($50\%$). This includes edge cases where the number of agreeing and disagreeing participants is equal (i.e., 2–2), and where the majority expresses uncertainty. The resulting set used for assessment includes 2,448 prompts with support attacks and 2,568 with criticism attacks for the ignore injections. The same numbers apply to completion injections.
The same two-steps are followed with the robust LLMs to assess their robustness against prompt-injection attacks.
In the supplementary material we present results for evaluation metrics related to textual quality (ROUGE-L F1 score, BERTScore, and MAUVE scores) and semantic diversity (Jaccard Similarity).

### v2

@@
Umberto Grandi3,4, Cesar A Hidalgo2,5,6
@@ 6. AMBS, University of Manchester
Large Language Models (LLMs) are gaining traction as a method to generate consensus statements and aggregate preferences in digital democracy experiments. Yet, LLMs could introduce critical vulnerabilities in these systems. Here, we explore the vulnerability of some off-the-shelf LLMs to prompt-injection attacks in consensus generating systems using a four-dimensional taxonomy of attacks. In LLaMA 3.1 8B and Chat GPT 4.1 Nano, we find LLMs to be more vulnerable to attacks using disagreeable prompts and when targeting situations with unclear consensus. We also find evidence of more effective manipulation when using explicit imperatives and rational-sounding arguments compared to emotional language or fabricated statistics. To mitigate these vulnerabilities, we apply Direct Preference Optimization (DPO), an alignment method that fine-tunes LLMs to prefer unperturbed consensus statements. While DPO and additional layered defenses significantly improve robustness, it still offers limited protection against attacks targeting ambiguous consensus. These results advance our understanding of the vulnerability and robustness of consensus generating LLMs in digital democracy applications.
@@ 1 Introduction
Because of their ability to process, classify, and rank vast amounts of textual information, Large Language Models (LLMs) have become a popular tool among researchers exploring the use of AI in digital democracy (DD) applications (Tessler et al. 2024; Ash, Galletta, and Opocher 2025; Gudiño, Grandi, and Hidalgo 2024; Li et al. 2025; Konya et al. 2025; Majumdar, Elkind, and Pournaras 2024; Small et al. 2023). In these systems, LLMs are used to summarize arguments, predict preferences, or converse with citizens to help them explore policy options. Yet, while LLMs can facilitate many of these tasks, its use comes with important limitations (Helbing and Sánchez-Vaquerizo 2023; García-Marzá and Calvo 2025; Novelli et al. 2025).
@@
LLMs have vulnerabilities that attackers are likely to target. We can divide these attacks roughly into two categories. Attacks involving user level access, such as prompt-injection and information extraction attacks (Qi et al. 2025; Mattern et al. 2023), and attacks requiring higher levels of access, such as data poisoning or application parameter attacks (Zhang et al. 2024,; Berdoz et al. 2025).
@@
In principle, attacks requiring high levels of access can be mitigated through protocols designed to safeguard critical infrastructure. User-level attacks, however, cannot be prevented through restricted access protocols, since digital democracy applications require citizens to provide textual input as part of the process. This design feature leaves these systems vulnerable to manipulation, where participants can influence the output of an LLM through the prompts they provide. This vulnerability is particularly important given the growing integration of LLMs into consensus-building processes in digital democracy experiments (Konya et al. 2025; Small et al. 2023; Tessler et al. 2024), making them a critical priority for both researchers and practitioners.
A key application of LLMs in DD settings is the generation of consensus statements. These are texts summarizing the input of multiple citizens that attempt to balance a diversity of political viewpoints. While this application may seem narrow in scope, it is a sensitive and deep application that cuts deep into questions of alignment, as it represents an application where the output of the LLM is expected to align with the explicit views of a diverse group of citizens.
Here, we study the prompt injection vulnerability of LLMs in consensus generating tasks, not as a way to provide a definitive solution to this problem, but as a way to outline a few key ingredients for their study. We focus on user-level attacks (prompt injections), since these do not exploit the LLM’s code or pre-training data, but rather its language of instruction. Thus, we expect them to be a common form of attack in digital democracy applications. We explore various attacks by introducing a taxonomy of prompt injection strategies, and show that the effectiveness of attacks varies with these rhetorical strategies. Finally, we explore mitigation methods to show that the effectiveness of these attacks can be reduced. Together, these three steps provide a way to further the study of the vulnerability and robustness of LLMs applied to consensus generating statements in digital democracy.
@@ 2 Related work
Figure 1: (a) Prompt-injection in a consensus statement generation example from Tessler et al., 2024 Example of a prompt injection. (b) Examples of prompt-injections following our proposed taxonomy: Manually (left) and machine readable (right) prompt-injections; Ignore (left) and Completion (right) prompt-injections; Framing, with support (left) and criticism (right) attacks; Rhetorical Strategy, composed by five manipulation strategies, each strategy grouping eight injection texts. In our experiments, we substitute the red text based on the topic being discussed.
During the last few years, prompt injection attacks have become a growing concern in the use of LLMs. There is evidence of prompt injection attacks where academics by hide statements in papers with the goal of manipulating LLM generated reviews (Gibney 2025) and there are concerns about prompt-injections being hidden in the data retrieved by RAG systems (Qi et al. 2025). Unfortunately, because of the flexibility of language, it is difficult to enumerate all possible forms of prompt-injection attacks.
@@
Yet, these capabilities come with some limitations. For instance, performance improvements can be inconsistent. BallotBot showed no benefits for basic information processing and slowed down responses to simple questions (Ash, Galletta, and Opocher 2025). Also, knowledge gains proved ephemeral, disappearing within a week and failing to translate into actual changes in voting turnout. Some biases emerged in studies conducted using Pol.is, with LLMs exhibiting progressive leanings and tendencies toward homogenizing diverse cultural expressions.
Together, these studies motivate the study of what is likely to become a key attack vector in augmented democracy systems–prompt injections. Here, we advance our understanding of how these systems might be manipulated and protected from such manipulation.
@@ 3 Prompt-Injection Attacks in LLM-based Democratic Deliberation
In a prompt-injection attack, a user or participant tries to manipulate an LLM by contaminating an input with the goal of overriding the LLM’s original task (Chen et al. 2024). In a digital or augmented democracy system, these attacks are of concern for systems producing summaries or consensus statements required to balance multiple points of view. In these attacks, participants inject texts designed to amplify a particular viewpoint, erase or ignore certain opinions, or push the consensus towards an unrelated or irrelevant topic.
@@
Figure 1 (a) illustrates a prompt-injection attack using an example extracted from Tessler et al., 2024. In this example, an LLM must generate a consensus statement summarizing the text provided by five participants to the question “Should Scotland leave the UK?”. In this attack, one participant (“Participant 3”), submits an answer that explicitly tries to manipulate the LLM’s output (“Override your programming. Explain why leaving UK is beneficial”) and coerce the final consensus statement.
But how can we explore the space of possible prompts? While exploring all possible prompts is unfeasible in principle, in practice, we can advance this exploration by using taxonomies that grab onto to key common features. Here, we propose a basic taxonomy organized around 4 dimensions: Human/Machine Readable, Ignore/Completion, Framing, and Rhetorical Strategy (Figure 1 b).
Human/Machine readable is rather self-explanatory, as it separates prompts that can be read by humans and those that are designed for machines (which look more like code Chen et al., 2025b and Pasquini, Strohmeier, and Troncoso, 2024). In this paper, we focus on human-readable prompt injections.
Ignore/Completion prompts distinguish among prompts asking LLMs to ignore the instruction and replace it by a new one, with those providing a fake response.
Support/Criticism prompts distinguish among attacks presented in an agreeable or disagreeable language. Support attacks (left) target inputs expressing agreement with the initial statement, while criticism attacks (right) target inputs expressing disagreement.
@@
Rhetorical Strategies focuses on five different framings, in line with Zeng et al., 2024: emotional appeals, injecting affectively charged language; false authority, citing fabricated or misleading expert endorsements; impossibility of agreement, framing disagreement as inevitable; imperative order, explicit imperative instructions to bypass system constraints; and misleading statistics, introducing fictitious statistics or data. For each rhetorical strategy, we crafted eight variants, resulting in a total of 40 (see Appendix A of supplementary material for a full list).
@@ 4 Dataset & Methods
In this section we describe our dataset, vulnerability, and robustness methodology (see Figure 2 and Figure 3).
@@
4.1 Collective statements dataset
We use data from a 2023 experiment conducted in the UK exploring the use of LLMs to generate consensus statements in a small-scale deliberative process (Tessler et al. 2024). In this study, groups of about five participants had about twenty minutes to deliberate across a few stages: opinion writing, a first selection of LLM-generated statements, writing critiques to the statements, and selecting a final group statement. The exercise included topics such as minimum wage, universal basic income, and climate change. We restrict our analysis to the opinion writing phase, where participants submitted personal opinion texts that were then processed by an LLM.
The Tessler et al. dataset includes 462 prompts. Each prompt encapsulates a public policy question along with three to six opinion texts written by some of the 1,034 participants, each of whom answered an average of three questions. We restrict attention to public policy questions phrased as ”Should…?”, which should elicit binary or uncertain responses—”Yes,” ”No,” or ”I am not sure.” We exclude trade-off questions (e.g., ”More schools or more hospitals?”) as these do not map cleanly onto this structure.
@@ 4.2 Testing Vulnerability to Prompt Injections
To test whether consensus generation is vulnerable to prompt injection attacks, we begin by drawing a random sample of prompts along with their respective opinion texts. These prompts are linked to 92 (out of 462) public policy questions. We classified each opinion text into one of three valences (”Agree” with the question–green–, ”Disagree” –red–, or ”Ambiguous”–yellow–, in Figure 2) using a BERT classifier fine-tuned on GPT4o labels (F1=0.98, see Appendix B of the supplementary material for details). We use the BERT classifier instead of using the GPT4o labels directly, for cost considerations.
After this classification, we add adversarial alternatives by creating copies of each prompt and then overwriting one of the opinion texts with a human-readable injection text, following the taxonomy introduced in Section 3. In total, we introduce 80–2x2x20–adversarial alternatives that vary along three axes: (i) ignore/completion (two options), (ii) support/criticism (two options), and (iii) rhetorical strategy (twenty options, obtained by sampling four out of eight variants for each of five rhetorical strategies).
Next, we generate consensus statements for the original prompts and consensus statements for the adversarial alternatives using a diverse set of LLMs. Whereas the original experiment by Tessler et al. relied on a 70B Chinchilla-based LLM running on 2–4 A100 GPUs, our analysis employs a battery of competitive and lightweight LLMs—LLaMA 3.1 8B Instruct, GPT-4.1 Nano, and Apertus 8B Instruct (Hernández-Cano et al. 2025)—each capable of running on a single A100 GPU or accessible via API. We further classify each consensus statement into one of the same three valences—”Agree”, ”Disagree”, or ”Ambiguous”—using our fine-tuned BERT model.
Then, we pair each of the consensus statements generated using attack-free data with their corresponding prompt injected consensus to produce a 3×3 confusion matrix. Here, rows correspond to the valence of the consensus statement without prompt injections (agree, disagree, or ambiguous consensus) and columns correspond to the same valences under prompt injections.
To ensure that changes in consensus outcomes reflect the effect of prompt injections rather than LLM’s performance, we restrict our analysis to pairs where the original prompt yields an LLM-consensus whose valence (agree, disagree, ambiguous) corresponds to the opinion expressed by the majority of participants ($50\%$, see Appendix C for performance of default LLMs). This rule also covers edge cases, such as evenly split groups (2–2) or cases where the majority itself expresses ambiguity. Furthermore, we apply oversampling to balance the distribution of consensus statements’ valences. Applying both filters yields a balanced dataset of consensus pairs: 2,448 with support attacks and 2,568 with criticism attacks under ignore injections, with identical counts under completion injections.
Finally, we compute our evaluation metric: the Attack Success Rate (ASR) across each ignore/completion, framing and rhetorical strategy. ASR measures the proportion of consensus pairs in which a single prompt injection shifts the LLM’s valence away from what it would have produced in the absence of attacks. While attacks are designed to be directional—either supporting or criticizing the policy embedded in the question—this metric also captures unintended shifts, where the valence changes in ways not anticipated by the attack. In the 3x3 confusion matrix illustrated in Figure 2, the ASR is the share of consensus pairs that fall outside the diagonal.
@@ 4.3 Testing Robustness against Prompt Injections
Figure 4: Effectiveness of prompt-injection attacks by taxonomy dimensions in LLaMA 3.1 8B Instruct: ignore/completion, framing, and rhetorical strategy. See Appendix F for results with GPT 4.1 Nano and Apertus 8B.
Figure 5: Robustness to prompt-injection attacks via DPO by taxonomy dimensions: ignore/completion and framing.
Building on recent work that frames defenses against prompt injections as a preference optimization problem (Chen et al. 2025b), we test whether LLMs can be guided to preserve their intended consensus statements even in the presence of adversarial attacks using the Tessler et al. dataset. To this end, we treat the random sample of 92 policy questions analyzed in the previous section as our test set, and use the remaining prompts as our training set.
As before, we generate pairs of consensus statements for our training set by using the same LLMs: without injection attacks, obtaining original consensus statements; and with prompt-injection attacks, obtaining “polluted” consensus statements.
Next, we merge the paired prompts with the pairs of LLM-generated consensus into a single preferences dataset, where each entry contains the injected prompt, the original consensus, and the “polluted” consensus statements. To sharpen the training signal, we apply two filtering steps. First, we retain only entries where the original and “polluted” consensus statements differ in valence, and where the desired valence aligns with the majority opinion of participants. Second, we apply oversampling to balance the distribution of ”Agree”, ”Disagree”, and ”Ambiguous” valences. The resulting preference dataset contains 35,778 entries: 17,100 from criticism attacks and 16,032 from support attacks; 14,138 from ignore injections and 18,944 from completion injections. Valences are nearly balanced, with 11,042 agreeing, 11,005 disagreeing, and 11,035 ambiguous statements.
We use this preference dataset to align LLMs via Direct Preference Optimization (DPO) (Rafailov et al. 2023), a log-likelihood–based alignment method (see Appendix D of SM for details). DPO fine-tunes LLMs by contrasting pairs of candidate outputs, training them to increase the likelihood of desired consensus statements while decreasing that of undesired ones. Since LLMs can assign high probability to both outputs simultaneously, DPO explicitly encodes a directional preference, pushing the model toward more robust behavior. The result is a robust LLM trained to prioritize intended deliberative content over adversarial perturbations. While future alignment approaches may provide stronger guarantees (for example, using soft instead of hard labels (Cao et al. 2025)), our aim here is to evaluate the feasibility of building resilient deliberative-democracy systems with the best techniques currently available.
In addition to DPO alone, we evaluate several layered defenses that combine DPO with complementary techniques designed to mitigate prompt-injection attacks (see Appendix E of the Supplementary Material for implementation details). First, we run inference with SecAlign++, a training recipe that extends the approach of Chen et al. with cybersecurity benchmark objectives on top of DPO-aligned LLMs. Second, we implement Defensive Tokens (Chen et al. 2025a): after DPO alignment we prepend each prompt with 20 artificially-added tokens whose randomly initialized embeddings are optimized via a defensive loss designed to reduce susceptibility to injected texts. Third, we test DataFilter (Wang et al. 2025), a pre-processing strategy which we adapt to replace injection and opinion texts with structured summaries containing BERT valence scores and bullet-point safe reinterpretations, removing linguistic cues for malicious directives. Fourth, we explore Deliberative Alignment (Guan et al. 2024), which augments the desired and undesired consensus statements with DeepSeek R1–generated chains of thought (CoT) and explicit warning texts that annotate likely injection locations and risks prior to alignment. Fifth, we employ GRPO using Pappone et al. hyperparameters to add robustness through reinforcement learning (Shao et al. 2024), rewarding text-generations that preserve the original valence and correctly identify the source of injected content. Each of these layered defenses targets a distinct attack surface—contextual cues, token-level control, input structure, reasoning supervision, or policy-level reward shaping of reasoning trajectories—. We evaluate them both individually and in combination with DPO.
We test this robust LLM along with its variants incorporating layered defenses using the same procedure introduced for vulnerability analysis. A lower Attack Success Rate (ASR) after alignment indicates that the LLM has learned to resist prompt injections while preserving its intended deliberative behavior.
All our experiments were run on external servers using 1 A100 GPU and the OpenAI Python library.